◆ ◆course contents » onmer□Notesà Evaluate-Feedback 습 Printe, Info » Lecture Aluminum Platinum The bimetallic strip above is made of aluminum (coefficient of liner expansion 24 x10-6 c1) and platinum (coefficient of liner expansion 9 x10-6 c-1) When this strip is held down at the left end and heated, it wil OA: bend downward B: bend upward C: remain straight, while stretching its length Suomit Answer Tries 0/20 This discussion is closed Send Feedback
Answer #1

Coefficient of thermal expansion is defined as : the change in length per unit change in temperature per unit length.

i.e.  \small \alpha = \frac{\Delta l}{l\Delta T}

Where, \small \alpha is coefficient of linear thermal expansion

\small \Delta l is change in length

  \small \Delta T is change in temperature

or, we can write that

\small \Delta l\propto \alpha

Given, \small \alpha _{Aluminium} > \alpha _{Platinum}

So,   \small \Delta l _{Aluminium} > \Delta l _{Platinum}

i.e. if we increase the temperature, the aluminium rod will expand more than platinum rod.

As given in the figure, the aluminium rod is in upper side and platinum is in lower side. So, when aluminium will expand more than platinum, the platinum rod will apply a force on the ends of aluminium rod. As a result the rod will bend downward.

For any doubt please comment and please give an up vote. Thank you.